# Multi-Objective Neural Network Assisted Design Optimization of Soft Fin-Ray Grippers for Enhanced Grasping Performance

### v1

@@ Abstract
Soft Fin-Ray grippers can perform delicate and careful manipulation, which has caused notable attention in different fields. These grippers can handle objects of various forms and sizes safely. The internal structure of the Fin-Ray finger plays a significant role in its adaptability and grasping performance. However, modeling the non-linear grasp force and deformation behaviors for design purposes is challenging. Moreover, when the Fin-Ray finger becomes more rigid and capable of exerting higher forces, it becomes less delicate in handling objects. The contrast between these two objectives gives rise to a multi-objective optimization problem. In this study, we employ finite element method (FEM) to estimate the deflections and contact forces of the Fin-Ray, grasping cylindrical objects. This dataset is then used to construct a multilayer perception (MLP) for prediction of the contact force and the tip displacement. The FEM dataset consists of three input and four target features. The three input features of the MLP and optimization design variables are the thickness of the front and supporting beams, the thickness of the cross beams, and the equal spacing between the cross beams. In addition, the target features are the maximum contact forces and maximum tip displacements in x- and y-directions. The magnitude of maximum contact force and magnitude of maximum tip displacement are the two objectives, showing the trade-off between force and delicate manipulation in soft Fin-Ray grippers. Furthermore, the optimized set of solutions are found using multi-objective optimal techniques. We use non-dominated sorting genetic algorithm (NSGA-II) method for this purpose. Our findings demonstrate that our methodologies can be used to improve the design and gripping performance of soft robotic grippers, helping us to choose a design not only for delicate grasping but also for high-force applications.
Index Terms:
@@
Soft Robotics, Optimization, Neural Networks, Finite Element Method, Fin-Ray Grippers
I Introduction
Robotic grippers used to be made of rigid parts and links in most cases, but recent progress in soft robotics and related fields such as material science has led to attract much attention to soft grippers, and boosting their progress. Soft grippers have the ability to handle a wider range of objects compared to their rigid counterparts, while also allowing for the use of simpler control frameworks[1]. The basic idea of the Fin-Ray grippers has inspired biologically when Leif Kniese observed a phenomenon in fish; by applying force to the structure of the fish’s fins, it bends in the opposite direction of the force. By following up on this observation and studying the structure in collaboration with Rudolf Bannasch, the Fin-Ray Effect was introduced[2]. Fin-Ray grippers can be used to grasp different delicate and sensitive objects such as eggs and fruits[3]. In a study, An et al. designed an optimized gripper based on the Fin-Ray effect with an integrated linkage mechanism to grasp and harvest tomatoes. Their design provided a balanced force distribution on tomatoes and also included a blade at the tip of Fin-Ray finger that automatically cuts the stems while harvesting[4]. FEM is widely considered an accurate method, capable of simulating and modeling linear and multiphysics problems. In the context of soft robotics, it provides the ability to simulate complex structures and nonlinear elastic materials, two essential aspects of understanding soft robotics. However, when dealing with these problems, FEM is computationally expensive and may face convergence difficulties. Combining data-driven methods such as Machine Learning (ML) techniques with FEM offers new opportunities in the modeling and development of soft robotic systems. This combination creates a powerful tool for optimizing soft robotics structures, materials, and control[5]. Crooks et al. designed a novel gripper inspired by the traditional Fin-Ray grippers. Their gripper was activated via a motor-tendon mechanism and made of soft and hard materials, all printable as a complete gripper. Hard parts included crossbeams, supports, and fingernails. They compared their optimized design with the common Fin-Ray grippers both experimentally and by simulation. As a result of their design, gripper’s tip and structure could move more and perform better gripping with a lesser amount of force and actuation compared to the common Fin-Ray grippers. Also, the contact area was greater, causing a more stable grip, and the gripper was able to handle a heavier load[6]. Hazrat Ali et al. did a study on material selection for Fin-Ray grippers. They simulated a problem of single gripper finger’s response to a static force in ANSYS Workbench. They wanted to choose the best material according to their design goals[7]. Shan et al. proposed a kinetostatic model for multi-crossbeam soft fingers to accurately estimate contact forces during object grasping. Their approach not only enabled efficient force prediction but also incorporated the influence of varying stiffness across different finger segments, providing valuable insights for optimizing the design and performance[8]. Elgeneidy et al. used FEM to optimize the structural design of adaptive soft Fin-Ray fingers with the layer jamming mechanism for variable stiffening. This mechanism enabled a low contact force at first, and a higher amount of force in the jammed position because of the angle of ribs. The chosen optimization parameters were rib thickness, the first rib’s angle and the angle increment between ribs. They achieved enhanced shape adaptation and force generation, reduced initial contact forces which is desirable for delicate grasping, and increases final contact forces[9]. Xu et al. introduced a compliant adaptive Fin-Ray gripper. They changed the traditional Fin-Ray gripper by putting rigid parts into it, making its force-deformation behavior linear. Then, they trained a neural network using data from FEM. This made it possible to calculate external force from the finger’s deformation. However, the accuracy of the estimated force dropped as the applied force point moved from the middle to the finger’s ends[10]. De Barrie et al. developed a neural network for real-time prediction of contact forces and stress maps in soft Fin-Ray grippers. They prepared a dataset using FEM simulations of the gripper and different grasped objects, varying in size, shape, height, and angle of approach. Their network demonstrated promising results and was able to predict cases that were not in the training data. However, there were limitations in the real-world applications, including that the network’s performance dropped as the camera angle increased[11]. Deng et al. attempted to find an optimal Fin-Ray finger to improve its grasping. They defined different objects by a superquadratic model and, by considering the winding number, they presented the grasping quality function. They used numerical simulations, and considered thickness and spacing distance as the optimization parameters. They figured out that in real-world applications, computing the grasp quality function is hard due to the lack of precise determination of deformations. Also, the winding number is a geometric property and does not represent any aspect of the control framework[12]. Yao et al. proposed a two dimensional kinetostatic model of a soft Fin-Ray finger. The model was able to calculate the total contact force and deformations, and it was found out that the results obtained within their model exhibited similar outputs and accuracy compared to those of the FEM simulation. However, friction was not considered in the model. Next, they used this model to optimize the finger structure[13]. Yao et al. investigated the internal crossbeam structures of Fin-Ray grippers. They considered four different layouts, including one without crossbeams and used FEM to enhance the adaptability of the Fin-Ray fingers. They found out that removing the internal structure enhanced the finger’s ability to conform to delicate items while minimizing applied force[14]. Although Fin-Ray fingers can take the shape of an object very well, they have low performance in holding heavy loads. Topology optimization is considered a way to improve the grasping ability and safety factor of Fin-Ray fingers. Lakshmi Srinivas et al. compared fingers with different internal layouts both with and without topology optimization. Comparing the best layouts of two groups, while the non-optimized finger offered the best wrapping ability, the optimal one provided a superior balance of wrapping ability, structural strength, and lightweight design, making it an efficient choice for high-load and versatile grasping tasks[15]. Ghanizadeh et al. estimated the contact forces in soft Fin-Ray grippers for open-loop and closed-loop control purposes using FEM and experimental validation. However, this study considers only one internal structure for the Fin-Ray finger which limits the generalization of the findings[16]. Wang et al. proposed a physics-informed neural network (PINN) to model a soft Fin-Ray structure, where the minimum potential energy was integrated into the loss function based on elasticity theory. They trained two models, with and without the data from a real Fin-Ray finger. Their experiments showed that the PINN without the real data and the FEM had nearly same accuracy, while the PINN with the real data stood out and improved the accuracy[17].
This paper presents a method for optimizing the internal structure of the Fin-Ray finger. In the second section, we use FEM to estimate the contact force and tip displacement for the cylinders. In the third section, we employ neural networks to predict our desired target features. In the fourth section, we use a multi-objective optimization technique to find the set of all Pareto efficient solutions. In the last section, we prepare a conclusive summary of the paper.
II Finite Element Method
@@
In this section, we examine the Fin-Ray finger contact forces and tip displacement during interaction with a 20 mm diameter cylinder. We use the data from this part to train the MLP model in the following phases of our study.
@@
We use FEM to estimate the Fin-Ray finger contact force and tip displacement for different internal structures. Static structural analysis is employed to analyze contact forces with the ANSYS software. The cylindrical object with a diameter of 20 mm contacts the Fin-Ray finger at the midpoint of the fronting beam. This cylinder is considered to be linearly elastic with the properties of the acrylonitrile butadiene styrene (ABS) material found in manufacturing technical specifications. Thermoplastic polyurethane (TPU) is widely employed in soft grippers due to its flexibility and resilience[18]. In this study, all simulated Fin-Ray fingers have the same material of TPU 95A, which the standard library of ANSYS materials does not include[11]. Hence, we consider Young’s modulus of 26 MPa and Poisson’s ratio of 0.48 as the mechanical properties of this material[14]. Afterwards, we define the grasp frictional contact region and additional essential constraints for the simulations, such as the fixed displacement constraints. We also use the default mechanically programmed controlled mesh method with an adaptive size function. It is assumed that the Fin-Ray finger experiences a 2.5 cm base displacement in the direction of the cylindrical object. In addition, we conduct a convergence study to make sure that our numerical results are independent of the mesh.
@@
We perform FEM simulations for Fin-ray fingers with different internal structures to determine the tip displacement and estimate the contact force for each case. To this end, we change three design parameters of the Fin-Ray finger internal structure: thickness of the front and supporting beams, the thickness of the crossing beams, and the equal spacing between each crossbeam. The thicknesses of the support and front beams are considered equal in this study. These parameters are vital because of their importance in influencing our target features. The range in which these parameters have changed is summarized in Table I. We simulate all possible structures within this feasible range of thicknesses. Other internal structural parameters, such as the angle of inclination of the crossbeams and the length of the finger base, are maintained unchanged. The front beam and crossbeams are parallel to the y and x axes just before contact, respectively. An illustration of the Fin-Ray finger before and after contact with the cylindrical object is shown in Fig. 1, where $D_{x}$ and $D_{y}$ are the tip displacement of the Fin-Ray finger in the x- and y-directions. Furthermore, the base length, the fronting and supporting beam lengths, and width of the Fin-Ray finger are assumed to be 35 mm, 100 mm, and 30 mm, respectively.
Figure 1: An illustration of the Fin-Ray finger in simulation
TABLE I: Values for each design parameter of Fin-Ray finger
@@
Thickness of front and supporting beams 1.5-4 mm 0.5 mm
Thickness of crossbeams 0.6-1.6 mm 0.2 mm
Spacing between each crossbeam 10-20 mm 2 mm
The total displacement contour for one of the internal structures is depicted in Fig. 2 as an example of the results. Parameters for this specific internal structure are the front and supporting beam with thickness of 2 mm, the crossing beams with the thickness of 0.6 mm, and each crossbeam equal spacing of 16 mm.
Figure 2: Total deformation contour for the specified Fin-ray finger in meters
After each FEM simulation, we record values of the maximum contact force and the maximum tip displacement in both x- and y- directions to form a dataset. This dataset is used to train the MLP model in the next phase of the investigation. We note that maximum values of contact force and maximum tip displacement are always achieved at the maximum base displacement of the Fin-Ray finger, which is 2.5 cm. The variation of the contact force with the base displacement is displayed in Fig. 3 for the internal structure with fronting and supporting beam thickness of 2 mm, the crossing beams thickness of 0.6 mm, and each crossbeam equal spacing of 16 mm. For this proposed example, we can see that the force-displacement behavior is almost linear for the first 1.5 cm of base displacement, and in the last 1 cm of base displacement this behavior is nonlinear as a result of the layer jamming effect [19].
@@
III MLP Model
In this section, we derive the MLP model to approximate the behavior of our complex system. This model will be employed later to determine the optimal internal structure of the Fin-Ray finger.
The Min-Max feature scaling is chosen here as the normalization method. We use Pearson correlation heatmap as a graphical tool that displays the correlation between our variables, derived by Eq. 1 where $r$ is the correlation coefficient, $x_{i}$ is the values of the variable $x$ in the sample, and $\bar{x_{i}}$ is the mean of $x$. Also, $y_{i}$ denotes values of the variable $y$ in the sample, and $\bar{y_{i}}$ is the mean of $y$[20]. This heatmap approach suggests that the correlations between each input and target feature are significant, which shows the suitability of the selected features. Other exploratory data analysis (EDA) techniques are used to reduce the effect of any outlier and address similar problems. Hence, we organize the data to ensure accuracy, consistency, and suitability for the task in a proper manner.

### v2

# Soft Robotics; Fin-Ray Fingers; Soft Grippers; Multi-objective Optimization; Neural Networks; Finite Element Method journal: Engineering Reports

@@ Abstract
The internal structure of the Fin-Ray fingers plays a significant role in their adaptability and grasping performance. However, modeling the grasp force and deformation behavior for design purposes is challenging. When the Fin-Ray finger becomes more rigid and capable of exerting higher forces, it becomes less delicate in handling objects. The contrast between these two gives rise to a multi-objective optimization problem. We employ the finite element method to estimate the deflections and contact forces of the Fin-Ray fingers grasping cylindrical objects, generating a dataset of 120 simulations. This dataset includes three input variables: the thickness of the front and support beams, the thickness of the crossbeams, and the equal spacing between the crossbeams, which are the design variables in the optimization. This dataset is then used to construct a multilayer perceptron (MLP) with four output neurons predicting the contact force and tip displacement in two directions. The magnitudes of maximum contact force and maximum tip displacement are two optimization objectives, showing the trade-off between force and delicate manipulation. The set of solutions is found using the non-dominated sorting genetic algorithm (NSGA-II). The results of the simulations demonstrate that the proposed methodology can be used to improve the design and grasping performance of soft grippers, aiding to choose a design not only for delicate grasping but also for high-force applications.
keywords:
@@
Soft Robotics; Fin-Ray Fingers; Soft Grippers; Multi-objective Optimization; Neural Networks; Finite Element Method
journal: Engineering Reports
[…]
[label1]organization=School of Mechanical Engineering, College of Engineering, University of Tehran, Tehran, Iran
1 Introduction
Robotic grippers used to be made of rigid parts and links in most cases, but recent progress in soft robotics and related fields has attracted much attention to soft grippers and boosted their development. Soft grippers have the ability to handle a wider range of objects compared to their rigid counterparts, while also allowing for the use of simpler control frameworks [16].
Fin-Ray grippers are a type of soft gripper, and their basic concept was biologically inspired by Leif Kniese’s observation of a phenomenon in fish. By applying force to the structure of the fish’s fins, it bends in the opposite direction of the force. By following up on this observation and studying the structure in collaboration with Rudolf Bannasch, the Fin-Ray effect was introduced [13]. Fin-Ray grippers can be used to grasp different delicate and sensitive objects such as eggs [17] and fruits [2].
A key factor influencing the performance of Fin-Ray grippers is the finger design, and several studies focused on introducing new designs or improving existing ones based on the desired task. In a study, an optimized gripper based on the Fin-Ray effect with an integrated linkage mechanism was designed to grasp and harvest tomatoes. Their design provided a balanced force distribution on tomatoes and also included a blade at the tip of the Fin-Ray finger that automatically cuts the stems while harvesting [2]. Crooks et al. designed a novel gripper inspired by the traditional Fin-Ray grippers. However, their gripper was activated via a motor-tendon mechanism and made of a combination of soft and hard materials, all printable as a complete gripper. Hard parts included crossbeams, supports, and fingernails. They compared their optimized design with the common Fin-Ray grippers both experimentally and by simulation. As a result of their design, the gripper’s tip and structure could move more and perform better gripping, showing 15% greater deformation compared to a regular Fin-Ray finger. In addition, the contact area was larger, resulting in a more stable grip, and the gripper was able to handle about 40% heavier loads [5].
Besides the Fin-Ray finger design, the material from which they are made plays an important role in their performance. Hence, a study was conducted on material selection for Fin-Ray grippers. They simulated the response of a single gripper finger to a static force in ANSYS Workbench. Their objective was to choose the best material according to their design goals [1].
To optimize the design of Fin-Ray fingers, several geometric factors can be considered, including rib (crossbeam) thickness, the crossbeam angles, and the angle increment or spacing distance between crossbeams. The optimization problem is carried out to achieve enhanced shape adaptation and force generation, reduced initial contact forces, or improved performance during the layer jamming phase [9, 8]. However, the optimization variables are not always limited to these, since the finger’s internal structure can also be directly studied. Topology optimization is considered a way to improve the grasping ability and safety factor of Fin-Ray fingers. Lakshmi Srinivas et al. compared fingers with different internal layouts both with and without topology optimization. Comparing the best layouts of two groups, while the non-optimized finger offered the best wrapping ability, the optimal one provided a superior balance of wrapping ability, structural strength, and lightweight design, making it an efficient choice for high-load and versatile grasping tasks [11]. Yao et al. investigated the internal crossbeam structures of Fin-Ray grippers. They considered four different layouts, including one without crossbeams and used FEM to enhance the adaptability of the Fin-Ray fingers. They found that removing the internal structure enhanced the finger’s ability to conform to delicate items while minimizing applied force [21].
FEM is widely considered an accurate method, capable of simulating and modeling linear and multiphysics problems. In the context of soft robotics, it provides the ability to simulate complex structures and nonlinear elastic materials, two essential aspects of understanding soft robotics. However, when dealing with these problems, FEM is computationally expensive and may face convergence difficulties. Combining data-driven methods such as Machine Learning (ML) techniques with FEM offers new opportunities in the modeling and development of soft robotic systems. This combination creates a powerful tool for optimizing soft robotics structures, material selection, and control strategies [12].
De Barrie et al. developed a neural network for real-time prediction of contact forces and stress maps in soft Fin-Ray grippers. They prepared a dataset using FEM simulations of the gripper that interacts with different grasped objects, varying in size, shape, height, and angle of approach. Their network demonstrated promising results and was able to predict cases that were not in the training data. However, there were limitations in the real-world applications, including that the network’s performance dropped as the camera angle increased [6]. Yao et al. proposed a two dimensional kinetostatic model of a soft Fin-Ray finger. The model was able to calculate the total contact force and deformations, and it was found out that the results obtained within their model exhibited similar outputs and accuracy compared to those of the FEM simulation. However, friction was not considered in the model. Next, they used this model to optimize the finger structure [22]. Also, in another study, a kinetostatic model was proposed for multi-crossbeam soft fingers to accurately estimate contact forces during object grasping. Their approach not only enabled efficient force prediction but also incorporated the influence of varying stiffness across different finger segments, providing valuable insights for optimizing the design and performance [15]. Ghanizadeh et al. estimated the contact forces in soft Fin-Ray grippers for grasping cylindrical objects using FEM and experimental validation. However, this study considered only one internal structure for the Fin-Ray finger which limits the generalization of the findings [10]. Xu et al. introduced a compliant adaptive Fin-Ray gripper. They changed the traditional Fin-Ray gripper by putting rigid parts into it, making its force-deformation behavior linear. Then, they trained a neural network using data from the finite element method (FEM). This made it possible to calculate external force from the finger’s deformation. However, the accuracy of the estimated force dropped as the applied force point moved from the middle to the finger’s ends [20]. Wang et al. proposed a physics-informed neural network (PINN) to model a soft Fin-Ray structure, where the minimum potential energy was integrated into the loss function based on elasticity theory. They trained two models, with and without the data from a real Fin-Ray finger. Their experiments showed that the PINN without the real data and the FEM had nearly the same accuracy, while the PINN with the real data stood out and improved the accuracy [18].
The importance of optimizing the Fin-Ray finger structure has been widely recognized. Since FEM simulations are computationally expensive, this study aims to employ a neural network that mimics the behavior of a Fin-Ray finger and is trained on a small dataset obtained from FEM simulations. In this work, we use a population size of 500 and 100 generations when applying NSGA-II in the optimization process. This would require about 50,000 FEM runs, which is prohibitively expensive. Instead, we train the MLP by running FEM simulations for only 120 Fin-Ray fingers, enabling fast and accurate evaluation. Next, the MLP is used in a multi-objective optimization problem to identify the Fin-Ray finger designs. While surrogate-based approaches have been explored in some studies on Fin-Ray grippers, applying a neural network as the main evaluation model for multi-objective optimization of internal structural parameters has not been studied, and this work contributes to filling that gap. This approach demonstrates the feasibility of achieving good performance with a limited dataset while significantly reducing computational cost and time. In the final section, we provide the conclusion, summarizing the key results of the article.
2 Finite Element Method
@@
In this section, we examine the Fin-Ray finger contact force and tip displacement during interaction with a 20 mm diameter cylinder. We use the data from this part to train the multilayer perceptron model in the following phases of our study.
@@
We use FEM to estimate the Fin-Ray finger contact force and tip displacement for different internal structures. Static structural analysis is employed to analyze contact forces with the ANSYS software. The cylindrical object with a diameter of 20 mm contacts the Fin-Ray finger at the midpoint of the front beam. This cylinder is considered to be linearly elastic with the properties of the acrylonitrile butadiene styrene (ABS) material found in manufacturing technical specifications. Thermoplastic polyurethane (TPU) is widely employed in soft grippers due to its flexibility and resilience [19]. In this study, all simulated Fin-Ray fingers have the same material of TPU 95A, which is not included in the standard ANSYS material library [6]. Hence, we consider Young’s modulus of 26 MPa and Poisson’s ratio of 0.48 as the mechanical properties of this material [21]. It is assumed that the Fin-Ray finger experiences a 2.5 cm base displacement in the direction of the cylindrical object. Afterwards, we define the grasp frictional contact region and additional essential constraints for the simulations, such as the fixed displacement constraints. The friction coefficient is assumed to be 0.25 and the contact is defined as ”frictional” in ANSYS simulations [21]. We employ the default ANSYS mechanically controlled meshing algorithm, resulting in an average element size of 5.88 mm. The default ANSYS convergence criteria for force and displacement are used. SOLID186 elements are applied to 3D solid regions, and CONTA174 elements are used to model contact and sliding interactions between soft deformable surfaces and 3D solids. The total element count varies among different Fin-Ray configurations. In addition, we conduct a convergence study to make sure that our numerical results are independent of the mesh. A mesh dependency study for one of the Fin-Ray structures with 1.5 mm front and support beam thicknesses, 0.6 mm crossbeam thickness, and 10 mm spacing between crossbeams is presented in Fig. 1. It can be seen that, with errors of approximately 0.57% and 1.06% in total displacement and total force relative to the finest mesh, the selected mesh size of 5.88 mm demonstrates mesh independence.
Figure 1: Mesh convergence study: (a) Total displacement, force and number of elements and (b) Relative changes with respect to the finest mesh values.
@@
We perform FEM simulations for Fin-ray fingers with different internal structures to determine the tip displacement and estimate the contact force for each case. To this end, we change three design parameters of the Fin-Ray finger internal structure: thickness of the front and support beams, the thickness of the crossbeams, and the equal spacing between each crossbeam. The thicknesses of the support and front beams are considered equal in this study. These parameters are vital because of their importance in influencing our target features. The range in which these parameters are varied is summarized in Table 1. This range can be selected differently depending on the application for which the Fin-Ray finger is used. We simulate all possible structures within this feasible range of thicknesses. As can be seen there are $6\times 5\times 4=120$ possible combinations and a total of 120 distinct structures is successfully simulated and included in the dataset. Other internal structural parameters, such as the angle of inclination of the crossbeams and the length of the finger base, are remained unchanged.
Table 1: Values for each design parameter of Fin-Ray finger
@@
Thickness of front and support beams 1.5-4 mm 0.5 mm
Thickness of crossbeams 0.8-1.6 mm 0.2 mm
Spacing between each crossbeam 10-16 mm 2 mm
The front beam and crossbeams are parallel to the y and x axes just before contact, respectively. An illustration of the Fin-Ray finger before and after contact with the cylindrical object is shown in Fig. 2, where $D_{x}$ and $D_{y}$ are the tip displacements of the Fin-Ray finger in the x- and y-directions. The crossbeams, front beam and support beams are also displayed in Fig. 2. Furthermore, the base length, the front and support beam lengths, and the width of the Fin-Ray finger are assumed to be 35 mm, 100 mm, and 30 mm, respectively.
Figure 2: A schematic of the Fin-Ray finger in simulation
The total displacement contours for one of the internal structures is depicted in Fig. 3 as an example of the results. Parameters for this specific internal structure are front and support beam with thickness of 2 mm, the crossbeams with the thickness of 0.8 mm, and each crossbeam equal spacing of 16 mm. In addition to the deformation contour shown in Fig. 3, we calculate the corresponding strains using the FEM simulations. As a result, the maximum strain in all of the scenarios is small. Therefore, TPU 95A’s behavior can be approximated to behave linearly elastic in that range [21, 14].
Figure 3: Total deformation contours for the specified Fin-ray finger in meters
After each FEM simulation, we record values of the maximum contact force and the maximum tip displacement in both x- and y- directions to form a dataset. This dataset is used to train the MLP model in the next phase of the investigation. We note that maximum values of contact force and maximum tip displacement are always achieved at the maximum base displacement of the Fin-Ray finger, which is 2.5 cm. The variation of the contact force with the base displacement is displayed in Fig. 4 for the internal structure with front and support beam thicknesses of 2 mm, the crossbeams thickness of 0.8 mm, and each crossbeam equal spacing of 16 mm. For this proposed example, we can see that the force-displacement behavior is almost linear for the first 1.5 cm of base displacement, and in the last 1 cm of base displacement this behavior is nonlinear due to geometric and contact nonlinearities.
@@
3 MLP Model
In this section, we derive the MLP model to approximate the behavior of our complex system. This model will be employed later to determine the internal structure of the Fin-Ray finger. The Min-Max feature scaling is chosen here as the normalization method. We use the Pearson correlation heatmap as a graphical tool that displays the correlation between our variables, given by Eq. 1.